OpenStudy (anonymous):

I need some guidance: solving the system. 4x-y+3z=-2 3x+5y-z=15 -2x+y+4z=14

OpenStudy (anonymous):

I know that I need to elimante y. So I took the 1st and 3rd equation and added them.

OpenStudy (anonymous):

4x-y+3z=-2 -2x+y+4z=14 ________________ 2x+7z=12

OpenStudy (anonymous):

Then I did 2 and 3rd equations: 3x+5y-z=15 3x+5y-z=15 -2x+y+4z=14 10x-5y-20z=-70 _________________ ____________________ 13x-21z=-55

OpenStudy (anonymous):

now use the two equations that you got 2x+7z=12 13x-21z=-55 to solve for one of the variables (x or z) and go on and solve each variable one by one

OpenStudy (anonymous):

Then I need to take my answers from the first and second and add them together. 2x+7z=12 6x+21=12 13x-21z=-55 13x-21z=-55 _________________ Then I get lost?

OpenStudy (anonymous):

its suppose to be 36 at the end of the equation... let me see.

OpenStudy (anonymous):

(-1,4,2)

OpenStudy (anonymous):

you mean, you solved for x and your answer is wrong? oO

OpenStudy (dumbcow):

looks good so far... 6x +21z = 36 13x-21z =-55 ________________ 19x = -19 x = -1

OpenStudy (anonymous):

yeah i think dumbcow got it just keep plug and solve-ing dumbcow...when i say your name i feel i'm like making fun of you...i'm sory :x

OpenStudy (dumbcow):

(-1,4,2) is correct @ chiaroscuro dont worry about, its an inside joke ;)

OpenStudy (anonymous):

Thanks Y'all!!!